Jill Hawkins, Chief Marketing Officer for Millennium Sporting Equipment (MSE), Inc., must decide whether to introduce a mid-priced version of the firm's Voyager line of high performance recreational bicycles next year. The new model (V-330) has a unit variable cost of $180, and will be sold to retailers at a price of $390. Projections made by the product development team indicate that the V-330 would achieve a sales volume of 50,000 units next year, its first year of commercialization. Sixty percent of the V-330’s first year's sales volume is expected to come from competitors' products and market growth, while the remaining 40 percent will come from two models currently sold by MSE. Sixty percent of this cannibalized volume will come from the higher-priced V-660 model, which has unit variable costs of $220 and is sold to retailers for $590. The remaining 40 percent will come from the economy-priced V-220 (priced at $250 to retailers with unit variable cost of $120). If V-330 is not introduced, sales of the existing models next year are expected to be 40,000 units for V-660 and 60,000 units for V-220. The fixed costs of launching V-330 have been forecast to be $3 million

. a) What is the cannibalization rate?

   b) How many units each of V-220 and V-660 will be cannibalized? (2 points)

c) How many incremental units of V-330 will be sold? (1 point)

d) Should Ms. Hawkins add the V-330 model to the product line? Why? (6 points)
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Millennium Sporting Equipment
a) What is the cannibalization rate?
Cannibalization rate is a marketing term. It means the % of new product sales which represents the sales lost of old product due to launch of new product.
Here it is mentioned that Sixty percent of the V-330’s first year's sales volume is expected to come from competitors' products and market growth, while the remaining 40 percent will come from two models currently sold by MSE.
So the cannibalization rate is 40%.
Answer b Amount $ Note
Sales volume of V-330            50,000.00 A
Cannibalization rate 40% B
Total Units to be cannibalized           20,000.00 C=A*B
60% of this will come from V-660            12,000.00 D=C*60%
40% of this will come from V-220              8,000.00 E=C*40%
So,
Units each of V-220 be cannibalized              8,000.00 See E
Units each of V-660 be cannibalized           12,000.00 See D
Answer c
Sales volume of V-330            50,000.00 See A
Less: Total Units to be cannibalized            20,000.00 See C
Incremental units of V-330 sold           30,000.00 F=A-C
Answer d
To decide to add the V-330 model to the product line or not we have to calculate incremental income from V-330 model. The workings are below:
Contribution lost of existing products V-660 V-220 Total Note
Units to be cannibalized            12,000.00            8,000.00 See D,E
Sell Price                 590.00               250.00 G
Variable cost per unit                 220.00               120.00 H
Contribution margin per unit                 370.00               130.00 I=G-H
Total Contribution lost      4,440,000.00 1,040,000.00 5,480,000.00 J=Units to be cannibalized*I
Incremental income statemnet-V-330 model:
V-330 Note
Sales volume            50,000.00 See A
Sell Price                 390.00 J
Variable cost per unit                 180.00 K
Contribution margin per unit                 210.00 L=J-K
Total Contribution 10,500,000.00 M=A*L
Less:
Fixed costs      3,000,000.00
Total Contribution lost      5,480,000.00 See J
Incremental income      2,020,000.00
Adding the V-330 model to the product line will have incremental income of $ 2,020,000. So yes the V-330 model should be added.